# Rubric Is All You Need: Enhancing LLM-based Code Evaluation With Question-Specific Rubrics

### v1

@@ Abstract.
Since the disruption in LLM technology brought about by the release of GPT-3 and ChatGPT, LLMs have shown remarkable promise in programming-related tasks. While code generation remains a popular field of research, code evaluation using LLMs remains a problem with no conclusive solution. In this paper, we focus on LLM-based code evaluation and attempt to fill in the existing gaps. We propose multi-agentic novel approaches using question-specific rubrics tailored to the problem statement, arguing that these perform better for logical assessment than the existing approaches that use question-agnostic rubrics. To address the lack of suitable evaluation datasets, we introduce two datasets: a Data Structures and Algorithms dataset containing 150 student submissions from a popular Data Structures and Algorithms practice website, and an Object Oriented Programming dataset comprising 80 student submissions from undergraduate computer science courses. In addition to using standard metrics (Spearman Correlation, Cohen’s Kappa), we additionally propose a new metric called as Leniency, which quantifies evaluation strictness relative to expert assessment. Our comprehensive analysis demonstrates that question-specific rubrics significantly enhance logical assessment of code in educational settings, providing better feedback aligned with instructional goals beyond mere syntactic correctness.
@@
*Equal Contribution
journalyear: 2025copyright: acmlicensedconference: 2025 ACM Conference on International Computing Education Research; August 3–6, 2025; Charlottesville, Virginia, United Statesprice: 15.00ccs: Social and professional topics Student assessmentccs: Computing methodologies Natural language processing
@@ 1. Introduction
The integration of LLMs into computing education has led to groundbreaking advancements, transforming both students and educators experiences (Hellas et al., 2024; Joshi et al., 2023; Hicke et al., 2023; Raihan et al., 2024). In computing education particularly, LLMs have demonstrated potential in generating instructional content (Sarsa et al., 2022; Phung et al., 2023), providing personalized tutoring and doubt solving(Liffiton et al., 2023; Kotalwar et al., 2025) as well as assisting with code evaluation and grading (Joshi et al., 2024; Alkafaween et al., 2024; Phung et al., 2023). Despite these advancements, there are significant gaps in using LLMs for code evaluation and grading. Most studies focus on generating feedback, helping students identify errors and improve their work (Alkafaween et al., 2024; Pankiewicz and Baker, 2023; Vaithilingam et al., 2022). Existing studies have demonstrated LLMs’ ability to provide meaningful insights into code quality, syntax, and logic (Pankiewicz and Baker, 2023; Anishka et al., 2024). Yet, grading, a critical and labor-intensive responsibility for instructors (Mandernach et al., 2013; Normann et al., 2023), has received comparatively little attention (Kiesler et al., 2023; Jacobs and Jaschke, 2024). The evaluation and grading of student code is a critical component of computing education, as it helps instructors assess students’ understanding of programming concepts and computational logic. This process consumes significant time and energy for instructors (Mandernach et al., 2013; Normann et al., 2023), making it a prime candidate for automation.
@@
”How effectively can an LLM-based grader evaluate student code using question-specific rubric compared to question-agonistic rubric? What measurable differences exist in evaluation quality and feedback specificity between these two approaches?”
@@
  • We present a new metric that measures the strictness or leniency of an evaluation system based on rubric. The proposed metric, Leniency, is relative and requires baseline data for comparison. (§5)
@@
By demonstrating the effectiveness of question-specific rubrics in LLM-based grading systems, we pave the way for more accurate and efficient automated evaluation tools. These tools have the potential to save instructors significant time and effort, allowing them to focus on other aspects of teaching and mentoring. Additionally, students stand to benefit from more detailed and contextually relevant feedback, which can help them identify areas for improvement and deepen their understanding of programming concepts.
@@ 2. Related Work
2.1. Limitations of Traditional Auto-grading Approaches
Conventional autograding systems have primarily relied on predefined test cases to evaluate student submissions. Lobb and Harlow(Lobb and Harlow, 2016) developed Coderunner, an early system that executed student code and compared outputs against expected results derived from instructor-defined test cases. While this approach allowed for basic assessment of functional correctness, it presented several limitations. Keuning et al. (Keuning et al., 2016) identified that these systems often lacked the ability to understand underlying semantic errors or provide nuanced explanations.
A significant limitation of test-case-based approaches is the burden placed on instructors to create comprehensive test suites manually. Keuning et al (Keuning et al., 2016) highlighted the challenges associated with generating test cases that effectively cover edge cases and potential error conditions. Furthermore, as pointed out by Vreda Pieterse (Pieterse, 2013); traditional systems often focus exclusively on functional correctness, neglecting aspects such as code style, efficiency, and adherence to best practices. This results in an incomplete assessment of student programming abilities and fails to promote holistic coding skills development.
Our approach addresses these challenges by employing instructor-defined rubrics that provide specific evaluation criteria, enabling more meaningful feedback beyond binary judgments while maintaining scalability. Though rubrics are also manually intensive, they have great upside in helping instructors evaluate students consistently and objectively (Chowdhury, 2018), while providing students with clear expectations and constructive feedback to identify their strengths and weaknesses (Otero Avila et al., 2019).
2.2. Emergence of LLMs in Programming Assessment
The emergence of Large Language Models (LLMs) marks a significant paradigm shift in autograding approaches. Recent studies by Denny et al. (Koutcheme et al., 2025) have outlined how LLMs can leverage their extensive training data to grasp nuances of code, identify common errors, and explain complex concepts in an accessible manner.
Phung et al. (Phung et al., 2023) investigated the possibility of LLMs mimicking the style and quality of feedback provided by human tutors, benchmarking ChatGPT(OpenAI, 2020) and GPT-4(OpenAI, 2023) against human instructors in programming education contexts. Their results showed that advanced models like GPT-4 can approach human-like assessment capabilities in specific scenarios, though significant gaps remain in others. Azaiz et al. (Azaiz et al., 2024) examined the effectiveness of GPT-4 in generating feedback, indicating its capability to provide insights beyond basic correctness checks.
Beyond simply evaluating code correctness, LLMs demonstrate potential for enhancing feedback quality. Leinonen et al. (Hellas et al., 2023) analyzed techniques for using LLMs to improve programming error messages, making them more informative and actionable for students. Kiesler et al. (Kiesler et al., 2023) explored how LLMs can provide formative feedback on programming assignments, suggesting that these models can offer more detailed and constructive guidance compared to traditional automated systems.
Though the above work highlights the potential of LLMs in the field, many studies also raised questions on the reliability and consistency of LLM-based evaluation (Seo et al., 2025) While the majority of the cited work focuses purely on providing effective feedback, we comprehensively also cover grading assignments which is essential for both instructors and students, especially for a large-scale programming course. Our approach using question-specific rubrics also ensures structured evaluation criteria that combine human expertise with AI capabilities to provide consistent, contextual feedback.
2.3. Major Frameworks and Approaches
Several frameworks have emerged to harness LLMs for programming assessment, and we are focusing on the ones mentioned in Table 1. CodeBERTScore (Zhou et al., 2023) marked a significant milestone in code evaluation by leveraging pre-trained BERT (Devlin et al., 2019) models to encode semantic vectors of reference and generated code, measuring the similarity between these vectors rather than relying on token-matching approaches. This method improved upon earlier techniques like traditional BLEU (Papineni et al., 2002) (which treated code as mere text) and its specialized derivative CodeBLEU (Ren et al., 2020), which had begun addressing code’s unique properties through weighted n-gram matching, AST comparison, and data-flow analysis. Despite CodeBERTScore’s innovations in capturing semantic meaning, it still faces a fundamental limitation: context similarity doesn’t necessarily represent semantic similarity, resulting in suboptimal performance when evaluating functionally identical code implemented with different approaches—a particular challenge in educational environments.
Recent work has explored using LLMs directly for code evaluation without relying on test cases. ICE-Score (Zhuo, 2024) pioneered this approach, making it more applicable in scenarios where test development is impractical. Despite its innovation, this approach demonstrates limited correlation with human judgment and remains susceptible to LLM hallucinations, particularly when evaluating complex code with intricate semantics.
Similarly, CodeJudge (Tong and Zhang, 2024) presents a promising approach leveraging ”slow thinking” to guide LLMs in evaluating code semantics. By decomposing evaluation into step-by-step analysis followed by summarization, and introducing a taxonomy of code inconsistencies with severity levels, it achieves better correlation with semantic correctness compared to existing methods. However, both approaches lack the context of specific programming assignments, limiting their effectiveness for educational assessment.
The techniques discussed above though have application in education are primarily used as benchmarks for code generation by LLMs for task description hence lack context and are unable to generate any meaningful feedback. Beyond these, we focused on rubric-based techniques for code evaluation and found a huge gap in methods exploring the same. For ease of our discussion, we further subdivide rubric-based programming evaluation on the basis of the Rubric specificity into a generic or question-agnostic rubric against a question-specific rubric accounting for the logical requirements from the problem.
While rubrics provide significant benefits for assessment, question-agnostic rubrics face challenges, particularly with automated grading systems for programming assignments. These general rubrics often fail to capture the nuances of specific programming problems, leading to misaligned evaluations, still struggling with grading feedback compared to human tutors (Phung et al., 2023). Their experiments showed that both GPT-4 and ChatGPT frequently misidentify code issues and incorrectly assign points for general correctness and edge cases when using general rubrics, suggesting question-specific rubrics might be more effective for automated assessment systems. More recent developments include domain-specific approaches like CodEV (Tseng et al., 2024), which leverage LLMs, Chain of Thought (CoT), LLM ensembles and a question agnostic rubric to improve score accuracy and consistency while agreement tests help ensure reliable feedback and precise code review comments.
Xie et al (Xie et al., 2024) propose a systematic approach to rethinking automated assessment for short answer with LLMs through their ”Grade Like a Human” system. This multi-agent system divides the grading process into three key stages: (1) developing context-aware grading rubrics that consider both the questions and student answers, (2) providing accurate scoring with customized feedback under these rubrics, and (3) conducting post-grading review to ensure assessment fairness and accuracy. It was designed primarily for short answer grading and hence not directly applicable to the programming assignment evaluation paradigm.
Building on previous methods, our approaches distinctly separate logical reasoning from syntax evaluation, offering practical guidelines for educational contexts where logical understanding outweighs syntactic correctness. Our multi-tiered rubric framework addresses programming education needs by simulating human grading practices, prioritizing conceptual understanding over minor implementation details. Our approaches also differentiate themselves by introducing a systematic rubric granularity framework— from complete to pointwise evaluation —allowing instructors to control assessment strictness based on pedagogical needs.
Table 1. Comparison of Code Assessment Techniques
@@ 2.3. Major Frameworks and Approaches
BLEU(Papineni et al., 2002)* $\times$ $\times$ $\times$ $\times$
CodeBLEU* (Ren et al., 2020) $\times$ $\times$ $\times$
CodeBERTScore*(Zhou et al., 2023)* $\times$ $\times$ $\times$
ICE-Score* (Zhuo, 2024) $\times$ $\times$ $\times$
CodeJudge* (Tong and Zhang, 2024) $\times$ $\times$
@@
*Focused on evaluating LLM-based code generation for task description
@@ 3. Dataset
We present two datasets, one based on OOP and second based on DSA. These courses are taken by all Computer Science undergraduates making them relevant to our study. Overall, our dataset contains 230 student submissions (80 from OOP and 150 from DSA). Full Dataset is available on Github  [^1] https://github.com/codeeval18/rubric-code-eval. Below, we provide the high level details about the dataset:
@@ 3.1. OOP Dataset Construction
For the OOP dataset, we rely on student submissions as part of a programming exam conducted in a OOP course at University X in fall 2024. The programming exam consisted of one Java programming question which further consisted of seven methods which needed to be implemented by the students. Students were provided with a structured starter template that included class declarations, helper functions, and predefined method scopes for implementation. A main function was also supplied to the students for testing and validating their solutions. The tasks in these methods focussed on applying object-oriented programming principles to handle file I/O, data filtering, and data updates.
  1. (1)
    Detailed Problem Statement: A detailed problem statement provided by the professor, outlining clear, step-by-step instructions for implementing each function in every part of the question.
  2. (2)
  3. (3)
  4. (4)
    Model Solution: A model solution prepared by the professor, which includes implementations of all steps while strictly adhering to the guidelines in the problem statement and rubric.
  5. (5)
    • Student code: The student’s submitted Java code.
@@ 7 Player p = new Player();
Figure 1. Sample Question from OOP dataset and corresponding student submission for the same
As an example, in the above student code, the line List¡Player¿ players = new List¡Player¿(); is incorrect, as List is an interface and cannot be instantiated directly (Step 1). However, all subsequent steps are correct. The final score, determined by human annotators, is 8 out of 9, with individual step-wise marks assigned as 0, 1, 1, 1, 1, 1, 1, 1, 1.
@@ 3.2. DSA Dataset Construction
To ensure diversity and comprehensiveness, we selected DSA problems from the Geeks for Geeks (GFG) practice website [7], spanning 9 topics and 3 difficulty levels (easy, medium, and hard). The distribution of problems across topics and difficulty levels is shown in Table 3 and Figure 3.
@@
For each problem, the dataset contains:
  1. (1)
    Problem Statement - The problem statements were sourced from the descriptions provided on the website for each question. These include textual description of the problem, input size constraints and example input-output pairs to illustrate expected behavior. A sample problem statment is shown in fig 5
  2. (2)
  3. (3)
  4. (4)
    Rubric - The rubric was designed by considering different approaches to solving the question. Marks were assigned to each step based on its relevance and significance across various possible solutions, a sample rubric is shown in fig 6
  5. (5)
    Feedback and Marking - Two human graders collaboratively evaluated six different solutions per problem, following the predefined rubric. They also provided detailed feedback for each implementation step as shown in fig 7. Both graders, as final-year students with extensive programming experience, ensured a thorough and consistent assessment process.
A sample comprising student solution is shown below:

### v2

# Rubric Is All You Need: Improving LLM-based Code Evaluation With Question-Specific Rubrics

@@ Abstract.
Since the emergence of Large Language Models (LLMs) popularized by the release of GPT-3 and ChatGPT, LLMs have shown remarkable promise in programming-related tasks. While code generation using LLMs has become a popular field of research, code evaluation using LLMs remains under-explored. In this paper, we focus on LLM-based code evaluation and attempt to fill in the existing gaps. We propose multi-agentic novel approaches using question-specific rubrics tailored to the problem statement, arguing that these perform better for logical assessment than the existing approaches that use question-agnostic rubrics. To address the lack of suitable evaluation datasets, we introduce two datasets: a Data Structures and Algorithms dataset containing 150 student submissions from a popular Data Structures and Algorithms practice website, and an Object Oriented Programming dataset comprising 80 student submissions from undergraduate computer science courses. In addition to using standard metrics (Spearman Correlation, Cohen’s Kappa), we additionally propose a new metric called as Leniency, which quantifies evaluation strictness relative to expert assessment. Our comprehensive analysis demonstrates that question-specific rubrics significantly enhance logical assessment of code in educational settings, providing better feedback aligned with instructional goals beyond mere syntactic correctness.
@@
copyright: noneccs: Social and professional topics Student assessmentccs: Computing methodologies Natural language processing
@@ 1. Introduction
The integration of Large Language Models (LLMs) into computing education has led to groundbreaking advancements, transforming both students and educators experiences (Hellas et al., 2024; Joshi et al., 2023; Hicke et al., 2023; Raihan et al., 2024) . In computing education particularly, LLMs have demonstrated potential in generating instructional content (Sarsa et al., 2022; Phung et al., 2023), providing personalized tutoring and doubt solving (Liffiton et al., 2023; Kotalwar et al., 2025) as well as assisting with code evaluation and grading (Joshi et al., 2024; Alkafaween et al., 2024; Phung et al., 2023). Despite these advancements, there are significant gaps in using LLMs for code evaluation and grading. Most studies focus on generating feedback, helping students identify errors and improve their work (Alkafaween et al., 2024; Pankiewicz and Baker, 2023; Vaithilingam et al., 2022). Existing studies have demonstrated LLMs’ ability to provide meaningful insights into code quality, syntax, and logic (Pankiewicz and Baker, 2023; Anishka et al., 2024). Yet, grading, a critical and labor-intensive responsibility for instructors (Mandernach et al., 2013; Normann et al., 2023), has received comparatively little attention (Kiesler et al., 2023; Jacobs and Jaschke, 2024). The evaluation and grading of student code is a critical component of computing education, as it helps instructors assess students’ understanding of programming concepts and computational logic. This process consumes significant time and energy for instructors (Mandernach et al., 2013; Normann et al., 2023), making it a prime candidate for automation.
@@
”How effectively can an LLM-based grader evaluate student code using a question-specific rubric compared to a question-agnostic rubric? What measurable differences exist in evaluation quality and feedback specificity between these two approaches?”
@@
  • We present a new metric that measures the strictness or leniency of an evaluation system based on rubric. (§5)
@@
By demonstrating the effectiveness of question-specific rubrics in LLM-based grading systems, we pave the way for more accurate and efficient automated evaluation tools. These tools have the potential to save instructors significant time and effort, allowing them to focus on other aspects of teaching and mentoring. Additionally, students stand to benefit from more detailed and contextually relevant feedback, which can help them identify areas for improvement and deepen their understanding of programming concepts. We have made our dataset and code publicly available on HuggingFace [^1] https://huggingface.co/datasets/BITS-Pilani-GRC/RubricEval and GitHub [^2] https://github.com/BITS-Pilani-GRC/Rubric-Grader respectively.
@@ 2. Related Work
BLEU (Papineni et al., 2002) $\times$ $\times$ $\times$ $\times$
CodeBLEU (Ren et al., 2020) $\times$ $\times$ $\times$
CodeBERTScore (Zhou et al., 2023) $\times$ $\times$ $\times$
ICE-Score (Zhuo, 2024) $\times$ $\times$ $\times$
CodeJudge (Tong and Zhang, 2024) $\times$ $\times$
@@
Table 1. Comparison of LLM-Based Techniques for Grading and Feedback
2.1. Limitations of Traditional Auto-grading Approaches
Conventional autograding systems have primarily relied on predefined test cases to evaluate student submissions. Lobb et al. (Lobb and Harlow, 2016) developed a system that executed student code and compared outputs against expected results derived from instructor-defined test cases. While this approach allowed for basic assessment of functional correctness, these systems often lack the ability to understand underlying semantic errors or provide nuanced explanations (Keuning et al., 2016).
Additional limitation of test-case-based approaches is the burden placed on instructors to create comprehensive test suites manually. Keuning et al (Keuning et al., 2016) highlighted the challenges associated with generating test cases that effectively cover edge cases and potential error conditions. Furthermore, traditional systems often focus exclusively on functional correctness, neglecting aspects such as code style, efficiency, and adherence to best practices (Pieterse, 2013). Finally, test suites require successful compilation where even minor syntactical errors can render logically correct code non-functional. This results in an incomplete assessment of student programming abilities and fails to promote holistic coding skills development.
Our approach addresses these challenges by employing instructor-defined rubrics that provide specific evaluation criteria. This enables more meaningful feedback beyond binary judgments while maintaining scalability. Though rubrics are also manually intensive, they have great upside in helping instructors evaluate students consistently and objectively (Chowdhury, 2018), while providing students with clear expectations and constructive feedback to identify their strengths and weaknesses (Otero Avila et al., 2019).
2.2. Emergence of LLMs in Programming Assessment
The emergence of Large Language Models (LLMs) marks a significant paradigm shift in autograding approaches. Recent studies by Denny et al. (Koutcheme et al., 2025) have outlined how LLMs can leverage their extensive training data to grasp nuances of code, identify common errors, and explain complex concepts in an accessible manner. Multiple studies (Phung et al., 2023; Azaiz et al., 2024; Kiesler et al., 2023) have shown that advanced LLMs such as GPT-4 (OpenAI, 2023) can approach human-level feedback quality and deliver formative insights beyond simple correctness checks. Furthermore, Leinonen et al. (Hellas et al., 2023) analyzed techniques for using LLMs to improve programming error messages, making them more informative and actionable for students.
Though the above work highlights the potential of LLMs in the field, many studies also raised questions on the reliability and consistency of LLM-based evaluation (Seo et al., 2025). While the majority of the existing work focuses purely on providing effective feedback, we comprehensively also cover grading the assignments which is essential for both instructors and students, especially for a large-scale programming course. Our approach using question-specific rubrics also ensures structured evaluation criteria that combine human expertise with AI capabilities to provide consistent, contextual feedback.
2.3. LLM-based Techniques for Grading and Feedback
Table 1 compares the major LLM-based approaches for programming assessment. CodeBERTScore (Zhou et al., 2023) leverages pre-trained BERT (Devlin et al., 2019) models to encode semantic vectors of reference and generated code, measuring the similarity between these vectors rather than relying on token-matching approaches. This method improved upon earlier techniques like traditional BLEU (Papineni et al., 2002) (which treated code as mere text) and its specialized derivative CodeBLEU (Ren et al., 2020) (which incorporated weighted n-gram matching, AST comparison, and data-flow analysis). Despite CodeBERTScore’s innovations in capturing semantic meaning, it still faces a fundamental limitation: context similarity doesn’t necessarily represent semantic similarity, resulting in suboptimal performance when evaluating functionally identical code implemented with different approaches which is a particular challenge in educational environments.
ICE-Score (Zhuo, 2024) has explored using LLMs directly for code evaluation without relying on test cases. However, this approach demonstrates limited correlation with human judgment and remains susceptible to LLM hallucinations, particularly when evaluating complex code with intricate semantics. Similarly, CodeJudge (Tong and Zhang, 2024) leverages ”slow thinking” to guide LLMs in evaluating code semantics. By decomposing evaluation into step-by-step analysis followed by summarization, and introducing a taxonomy of code inconsistencies with severity levels, it achieves better correlation with semantic correctness compared to other methods. The above mentioned techniques are primarily for benchmarking the correctness and quality of code generated by LLMs but are unable to generate any meaningful feedback for student code submissions.
Beyond these approaches, we examined rubric-based techniques for code evaluation and identified a significant gap in research exploring such methods. Although rubric creation is as time-intensive for instructors as building test suites, the resulting rubrics provide far greater educational value to students. Unlike test suites, which provide only error-based feedback and do not work when the student code does not compile, rubrics deliver direct, structured feedback that systematically supports both learning and assessment. For ease of our discussion, we further subdivide rubric-based programming evaluation on the basis of rubric specificity: (1) question-agnostic rubric and (2) question-specific rubric. These are formally defined in Figure 1.
{definitionframe}
Rubric Types for Programming Problems: For a programming problem $P$ with description $D$ and expected solutions $S$), we define:
  • A question-agnostic rubric $R_{\text{agnostic}}$ as a set of evaluation criteria independent of $specific-logical$ $requirements$. Formally, $R_{\text{agnostic}}(P_{1})=R_{\text{agnostic}}(P_{2})$ for any distinct problems $P_{1},P_{2}$. These criteria often cover general aspects like code style, basic correctness, and efficiency.
  • A question-specific rubric $R_{\text{specific}}(P)$ as a set of evaluation criteria tailored to $unique$ $logical-requirements$ and $constraints$. Formally, $R_{\text{specific}}(P_{1})\neq R_{\text{specific}}(P_{2})$ for distinct problems $P_{1},P_{2}$, with criteria directly referencing $D$ and $S$.
Figure 1. Formal definitions of rubric types used in LLM-based code evaluation.
LLM-based techniques using question-agnostic rubrics often fail to capture the nuances of specific programming problems, leading to misaligned evaluations, and still struggle with grading feedback compared to human tutors (Phung et al., 2023). Phung et al. (Phung et al., 2023) showed that both GPT-4 and ChatGPT frequently misidentify code issues and incorrectly assign points for general correctness and edge cases when using general rubrics, suggesting question-specific rubrics might be more effective for automated assessment systems. More recent developments include domain-specific approaches like CodEV (Tseng et al., 2024), which leverage LLMs, Chain of Thought, LLM ensembles, and a question-agnostic rubric to improve score accuracy and consistency.
Xie et al. (Xie et al., 2024) propose a multi-agent “Grade Like a Human” system which creates context-aware rubrics for short-answer questions. It then uses these rubrics to score responses, give tailored feedback, and run accuracy checks. Because it targets short-answer grading, this approach does not transfer directly to programming assessments.
To the best of our knowledge, this study is the first to utilize question-specific rubrics for grading and providing feedback on student code submissions. Our method distinctly separates the assessment of logical reasoning from syntactic correctness, mirroring real-world practice where conceptual understanding is given preference over code syntax.
@@ 3. Dataset
We present two datasets, one based on Object Oriented Programming (OOP) and second based on Data Structures and Algorithms (DSA). These courses are taken by all Computer Science undergraduates making them relevant to our study. Overall, our dataset contains 230 student submissions (80 from OOP and 150 from DSA). Full Dataset is available on HuggingFace  [^3] https://huggingface.co/datasets/BITS-Pilani-GRC/RubricEval. Below, we provide the high level details about the dataset:
@@ 3.1. OOP Dataset Construction
For the OOP dataset, we rely on student submissions as part of a programming exam conducted in an OOP course at BITS Pilani in Fall 2024. The programming exam consisted of one Java programming question which further consisted of seven methods which needed to be implemented by the students. Students were provided with a structured starter template that included class declarations, helper functions, and predefined method scopes for implementation. A main function was also supplied to the students for testing and validating their solutions. The tasks in these methods focussed on applying object-oriented programming principles to handle file I/O, data filtering, and data updates.
@@ 7 Player p = new Player();
Figure 2. Sample Question from OOP dataset and corresponding student submission
  1. (1)
    Detailed Problem Statement: A detailed problem statement provided by the instructor, outlining clear, step-by-step instructions for implementing each function in every part of the question.
  2. (2)
  3. (3)
  4. (4)
    Model Solution: A model solution prepared by the instructor, which includes implementations of all steps while strictly adhering to the guidelines in the problem statement and rubric.
  5. (5)
    • Student code: The student’s submitted Java code.
For example, in the student code in Figure 2, in line 2, List¡Player¿ players = new List¡Player¿(); is incorrect, as List is an interface and cannot be instantiated directly (Step 1). However, all subsequent steps are correct. The final score, determined by human annotators, is 8 out of 9, with individual step-wise marks assigned as 0, 1, 1, 1, 1, 1, 1, 1, 1.
@@ 3.2. DSA Dataset Construction
To ensure diversity and comprehensiveness, we selected DSA problems from the Geeks for Geeks (GFG) practice website [7], spanning 9 topics and 3 difficulty levels (easy, medium, and hard). The distribution of problems across topics and difficulty levels is shown in Figure 4 and Figure 4. For each problem, the dataset contains:
  1. (1)
    Problem Statement - The problem statements were sourced from the descriptions provided on the website for each question. These include textual description of the problem, input size constraints and example input-output pairs to illustrate expected behavior.
  2. (2)
  3. (3)
  4. (4)
    Rubric - The rubric was designed by considering different approaches to solving the question. Marks were assigned to each step based on its relevance and significance across various possible solutions.
  5. (5)
    Feedback and Marking - Two human graders collaboratively evaluated six different solutions per problem, following the predefined rubric. They also provided detailed feedback for each implementation step. Both graders, as final-year students with extensive programming experience, ensured a thorough and consistent assessment process.
@@
A sample comprising student solution, problem statement, rubric, and grade & feedback is shared in Figure 5, Figure 6, Figure 7, and Figure 8 respectively.